What is th rotational inertia 1 of a 230N cylinder whose diameter is 406mm
Answer #1

The moment of inertia of a solid cylinder about its central axis is given by

I = \frac{MR^{2}}{2}

Where M = Mass = weight /g = 230/9.8 = 23.5 Kg

R = Radius = diameter /2 = 406mm/2 = 203 mm = 0.203 m

23.5 × 0.203. 0.48Kg.rn